• January 30th: An unannounced fibre maintenance in Halmstad caused an outage oh HH's green uplink to SUNET.
  • January 24th: A faulty card on transmission equipment at KTH caused multiple outages.
  • January 23rd: There was an outage in the link towards VR caused by a of routing protocol failure.
  • January 19th: A transmission node restart caused an outage on one of the SUNET core links between Kalix and Stockholm.
  • January 18th: There was a short flap on the south green network caused by human error during another intervention at our site in Växjo.
  • January 17th: There was an outage on the red link between SUNET and UMU caused by a damaged fiber.
  • January 16th: There was an outage between SUNET and TDC in Stockholm.
  • January 10th: There was a routing problem for SMHI caused by a configuration mistake.
  • January 3rd: There was an outage on the south red network affecting the redundancy to BTH, HKR, KAU, LU and MAH. The cause was miscommunication between our supplier and their contractor who measured two fibers instead of just the one which had been scheduled.
  • December 26th: Router issues caused an outage on MDH's green network.
  • December 21st: Interfaces restarted on one of the SUNET routers at BTH, causing a short outage on their red uplink.
  • December 21st: Fibre issues in Göteborg caused a short outage on CTH's green network.


  • January 31st: Service affecting work was performed on the transmission network in order to restore management to a node in Sundsvall.
  • January 29th: Our supplier performed reconnection of circuits in Göteborg.
  • January 25th: A planned intervention was performed to fix the issues following a fiber break in Vindeln.
  • January 25th: proj.sunet.se has been moved to another server.
  • January 23rd: A planned maintenance was performed to replace a broken FPC in one of the routers in Stockholm.
  • January 20th: The green edge router at MDH was upgraded to a newer software version.
  • January 19th: The SiteVision CMS was updated to the latest version.
  • January 18th: The optical equipment at the site in Hakkas was moved to a new location.
  • January 12th: Our supplier performed a planned measurement on a fiber between Kalmar and Bergskvara.
  • January 11th: Our provider was supposed to perform fiber measurement on the south green network that they failed to do but they found and replaced a bad patch which was causing a degraded signal in that fiber stretch.
  • January 3rd: There was a maintenance aimed at fixing a degraded signal issue. The maintenance was unsuccessful.






1 pcs 00:02

3 pcs 01:12

4 pcs 01:14


7 pcs 1d 01:26

5 pcs 14:46

12 pcs 1d 16:12


4 pcs 1d 07:14

1 pcs 00:10

5 pcs 1d 07:24

Other: DNSsec

0 pcs

0 pcs

0 pcs


0 pcs

0 pcs

0 pcs


0 pcs

2 pcs 00:18

2 pcs 00:18

Total onåbarhet för redundant anslutna kunder: 1180 minuter.

  • No labels